### v1

# Self-Supervised Learning for Time Series Analysis: Taxonomy, Progress, and Prospects

Abstract
Self-supervised learning (SSL) has recently achieved impressive performance on various time series tasks. The most prominent advantage of SSL is that it reduces the dependence on labeled data. Based on the pre-training and fine-tuning strategy, even a small amount of labeled data can achieve high performance. Compared with many published self-supervised surveys on computer vision and natural language processing, a comprehensive survey for time series SSL is still missing. To fill this gap, we review current state-of-the-art SSL methods for time series data in this article. To this end, we first comprehensively review existing surveys related to SSL and time series, and then provide a new taxonomy of existing time series SSL methods. We summarize these methods into three categories: generative-based, contrastive-based, and adversarial-based. All methods can be further divided into ten subcategories. To facilitate the experiments and validation of time series SSL methods, we also summarize datasets commonly used in time series forecasting, classification, anomaly detection, and clustering tasks. Finally, we present the future directions of SSL for time series analysis.
Index Terms:
Time series analysis, self-supervised learning, representation learning, deep learning
1 Introduction
Time series data abound in many real-world scenarios [1, 2], including human activity recognition [3], industrial fault diagnosis [4], smart building management [5], and healthcare [6]. The key to most tasks based on time series analysis is to extract useful and informative features. In recent years, Deep Learning (DL) has shown impressive performance in extracting hidden patterns and features of the data. Generally, the availability of sufficiently large labeled data is one of the critical factors for a reliable DL-based feature extraction model, usually referred to as supervised learning. Unfortunately, this requirement is difficult to meet in some practical scenarios, particularly for time series data, where obtaining labeled data is a time-consuming process. As an alternative, Self-Supervised Learning (SSL) has garnered increasing attention for its label-efficiency and generalization ability, and consequently, many latest time series modeling methods have been following this learning paradigm.
SSL is a subset of unsupervised learning that utilizes pretext tasks to derive supervision signals from unlabeled data. These pretext tasks are self-generated challenges that the model solves to learn from the data, thereby creating valuable representations for downstream tasks. SSL does not require additional manually labeled data because the supervisory signal is derived from the data itself. With the help of well-designed pretext tasks, SSL has recently achieved great success in the domains of Computer Vision (CV) [7, 8, 9, 10] and Natural Language Processing (NLP) [11, 12].
Figure 1: The proposed taxonomy of SSL for time series data.
With the great success of SSL in CV and NLP, it is appealing to extend SSL to time series data. However, transferring the pretext tasks designed for CV/NLP directly to time series data is non-trivial, and often fails to work in many scenarios. Here we highlight some typical challenges that arise when applying SSL to time series data.. First, time series data exhibit unique properties such as seasonality, trend, and frequency domain information [13, 14, 15]. Since most pretext tasks designed for image or language data do not consider these semantics related to time series data, they cannot be directly adopted. Second, some techniques commonly used in SSL, such as data augmentation, need to be specially designed for time series data. For example, rotation and crop are the commonly used augmentation techniques for image data [16]. However, these two techniques may break the temporal dependency of the series data. Third, most time series data contain multiple dimensions, i.e., multivariate time series. However, useful information usually only exists in a few dimensions, making it difficult to extract useful information in time series using SSL methods for other data types.
To the best of our knowledge, there has yet to be a comprehensive and systematic review of SSL for time series data, in contrast to the extensive literature on SSL for CV or NLP [17, 18]. The surveys proposed by Eldele et al. [19] and Deldari et al. [20] are partly similar to our work. However, these two reviews only discuss a small part of self-supervised contrastive learning (SSCL), which requires a more comprehensive literature review. Furthermore, a summary of benchmark time series datasets needs to be included, and the potential research directions for time series SSL are also scarce.
This article provides a review of current state-of-the-art SSL methods for time series data. We begin by summarizing recent reviews on SSL and time series data and then propose a new taxonomy from three perspectives: generative-based, contrastive-based, and adversarial-based. The taxonomy is similar to the one proposed by Liu et al. [21] but specifically concentrated on time series data. For generative-based methods, we describe three frameworks: autoregressive-based forecasting, auto-encoder-based reconstruction, and diffusion-based generation. For contrastive-based methods, we divide the existing work into five categories based on how positive and negative samples are generated, including sampling contrast, prediction contrast, augmentation contrast, prototype contrast, and expert knowledge contrast. For each category, we analyze its insights and limitations. Then we sort out and summarize the adversarial-based methods based on two target tasks: time series generation/imputation and auxiliary representation enhancement. The proposed taxonomy is shown in Fig. 1. We conclude this work by discussing possible future directions for time series SSL, including selection and combination of data augmentation, selection of positive and negative samples in SSCL, the inductive bias for time series SSL, theoretical analysis of SSCL, adversarial attacks and robust analysis on time series, time series domain adaption, pretraining and large models for time series, time series SSL in collaborative systems, and benchmark evaluation for time series SSL.
Our main contributions are summarized as follows.
  • New taxonomy and comprehensive review. We provide a new taxonomy and a detailed and up-to-date review of time series SSL. We divide existing methods into ten categories, and for each category, we describe the basic frameworks, mathematical expression, fine-grained classification, and detailed comparison. To the best of our knowledge, this is the first work to comprehensively and systematically review the existing studies of SSL for time series data.
  • Collection of applications and datasets. We collect resources on time series SSL, including applications and datasets, and investigate related data sources, characteristics, and corresponding works.
  • Abundant future directions. We point out some key problems in this field from both applicative and methodology perspectives, analyze their causes and possible solutions, and discuss future research directions for time series SSL. We strongly believe that our efforts will ignite further research interests in time series SSL.
The rest of the article is organized as follows. Section 2 provides some review literature on SSL and time series data. Section 3 to Section 5 describe the generation-based, contrastive-based, and adversarial-based methods, respectively. Section 6 lists some commonly used time series data sets from the application perspective. Section 7 discusses several promising directions of time series SSL, and Section 8 concludes the article.
2 Related Surveys
In this section, several recent reviews on SSL and time series analysis are scrutinized.
2.1 Surveys on SSL
The surveys on SSL can be categorized by different criteria. In this paper, we outline four widely used criteria: learning paradigms, pretext tasks, components/modules, and downstream tasks. The details are described as follows.
2.1.1 Learning paradigms
This category focuses on model architectures and training objectives. The SSL methods can be roughly divided into the following categories: generative-based, contrastive-based, and adversarial-based methods. The model architectures of these categories are shown in Fig. 2 and can be summarized as:
  • The generative-based approach first uses an encoder to map the input $\mathbf{x}$ to the representation $\mathbf{z}$, and then a decoder to reconstruct $\mathbf{x}$ from $\mathbf{z}$. The training objective is to minimize the reconstruction error between the input $\mathbf{x}$ and the reconstructed input $\hat{\mathbf{x}}$.
  • The contrastive-based approach is one of the most widely used SSL strategies, which constructs positive and negative samples through data augmentation or context sampling. The model is then trained by maximizing the Mutual Information (MI) between the two positive samples. Contrastive-based methods usually use a contrastive similarity metric, such as InfoNCE loss [22].
  • The adversarial-based approach usually consists of a generator and a discriminator. The generator generates fake samples, and the discriminator is used to distinguish them from real samples.
Using the learning paradigm as a taxonomy is arguably the most popular among the existing SSL surveys, including [23, 20, 24, 25, 26, 27, 28]. However, not all surveys cover the above three categories. The readers are referred to these surveys for more details. In Table I, we also provide the data modalities involved in each survey, which can help readers quickly find the research work closely related to them.
Figure 2: Learning paradigms of SSL.
2.1.2 Pretext tasks
The pretext task serves as a means to learn informative representations for downstream tasks. Ericsson et al. [29] provides a very comprehensive review of pretext tasks for multiple modalities, including image, video, text, audio, time series, and graph. The various self-supervised pretexts are divided into five broad families: transformation prediction, masked prediction, instance discrimination, clustering, and contrastive instance discrimination. Unlike the learning-paradigm-based criterion, the pretext-task-based criterion is also related to data modality. For example, Jing and Tian [18] summarize the self-supervised feature learning methods on image and video data, and four categories are discussed: generation-based, context-based, free semantic label-based, and cross modal-based, where cross-modal-based methods construct learning task using RGB frame sequence an optical flow sequence, which are unique features in the video.
TABLE I: An overview of recent SSL surveys. The yellow part indicates the proportion of this modality in all six modalities. It can be seen that this article is mainly concerned with SSL on time series data.
Paper Modality
Image Video Audio Graph Text Time Series
[21] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[23] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[29] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[30] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[17] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[20] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[18] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[31] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[24] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[25] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[26] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[27] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[32] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[28] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ours [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2.1.3 Components and modules
The literature categorizing SSL methods according to their modules and components throughout the pipeline is also an important direction. Jaiswal et al. [17] and Le-Khac et al. [30] sort out the modules and components required in SSL from different perspectives. The pipeline can be divided into four components: positive and negative samples, pretext task, model architecture, and training loss. The basic intuition behind SSCL is to pull positive samples closer and push negative samples away. Therefore, the first step is to construct true positive and negative samples using data transformation or sampling techniques, such as data augmentation and context sampling on time series. It is worth noting that the choice of data augmentation techniques depends on the data modality. For example, the augmentation methods of image data include crop, color distort, resize, and rotate [16], while for time series, the augmentation method involves noise injection, window slicing, and window wrapping [33]. Furthermore, the sample generation method based on context sampling is mostly used for data with temporal dependencies. The second component is pretext tasks, which is a self-supervised task that acts as an important strategy to learn data representations using pseudo-labels [17]. The third component is model architecture, which determines how positive and negative samples are encoded during training. The major categories include end-to-end [16], memory bank [34], momentum encoder [35], and clustering [36].
2.1.4 Downstream tasks
The last way to categorize SSL literature is by downstream tasks. On the topic of speech processing, Latif et al. [31] summarize the development and challenges of self-supervised techniques in three different downstream tasks, namely automatic speech recognition, speaker recognition, and speaker emotion recognition, respectively.
2.2 Surveys on time series data
The surveys on time series data can be roughly divided into two categories. The first category focuses on different tasks, such as classification [37, 38], forecasting [39, 40, 41, 42], and anomaly detection [43, 44]. These surveys comprehensively sort out the existing methods for each task. The second category focuses on the key components of time series modeling based on deep neural networks, such as data augmentation [33, 45, 46] and model structure [47, 48]. [33] proposed a new taxonomy that divides the existing data augmentation techniques into basic and advanced approaches. [45] also provides a taxonomy and outlines four families: transformation-based methods, pattern mixing, generative models, and decomposition methods. Moreover, both [33] and [45] empirically compare different data augmentation methods for time series classification tasks. [47] systematically reviews transformer schemes for time series modeling from two perspectives: network structure and applications. The taxonomy proposed by Eldele et al. [19] and Deldari et al. [20] is somewhat similar to our proposed taxonomy, i.e., both taxonomies involve time series self-supervised contrastive learning methods. However, our taxonomy provides more detailed categories and more literature in the contrastive-based approach. More importantly, we also thoroughly review a large set of literature for the generative-based and adversarial-based approaches.
3 Generative-based Methods
In this category, the pretext task is to generate the expected data based on a given view of the data. In the context of time series modeling, the commonly used pretext tasks include using the past series to forecast the future windows or specific time stamps, using the encoder and decoder to reconstruct the input, and forecasting the unseen part of the masked time series. This section sorts out the existing self-supervised representation learning methods in time series modeling from the perspectives of autoregressive-based forecasting, autoencoder-based reconstruction, and diffusion-based generation. It should be noted that autoencoder-based reconstruction task is also viewed as an unsupervised framework. In the context of SSL, we mainly use the reconstruction task as a pretext task, and the final goal is to obtain the representations through autoencoder models.
3.1 Autoregressive-based forecasting
Given the current time-stamp $t$, the goal of an autoregressive-based forecasting (ARF) task is to forecast a window of length $K$ after time-stamp $t$ using the series before $t$, which can be expressed as:
$\hat{X}_{[t+1:t+K]}=f(X_{[1:t]}),$ (1)
where $\hat{X}_{[t+1:t+K]}$ represents the target window, and $K$ represents the length of the target window. When $K=1$, (1) is a single-step forecasting model, and it is a multi-step forecasting model when $K>1$. $X_{[1:t]}$ represents the input series before time $t$ (including $t$), which is usually used as the input of the model. $f(\cdot)$ represents the forecasting model. The learning objective is to minimize the distance between the predicted target window and the ground truth, thus the loss function can be defined as:
$\mathcal{L}=D(\hat{X}_{[t+1:t+K]},X_{[t+1:t+K]}),$ (2)
where $D(\cdot)$ represents the distance between the predicted future window $\hat{X}_{[t+1:t+K]}$ and the ground-truth future window $X_{[t+1:t+K]}$, usually measured by the mean square error (MSE), i.e.,
$\mathcal{L}=\frac{1}{K}\sum_{k=1}^{K}(\hat{X}_{[t+k]}-X_{[t+k]})^{2}.$ (3)
In the time series modeling with autoregressive-based forecasting task as a pretext task, Recurrent Neural Networks (RNNs) are widely used thanks to their strong capability in spatiotemporal dynamic behavior modeling or sequence prediction [41, 42, 49, 50]. Therefore, it is also naturally applied in the pretext task based on autoregressive forecasting. THOC [51] constructs a self-supervised pretext task for multi-resolution single-step forecasting called Temporal Self-Supervision (TSS). TSS takes the L-layer dilated-RNN with Skip-Connection structure as the model. By setting skip length, it can ensure that the forecasting tasks can be performed with different resolutions at the same time. In addition to RNNs, the forecasting models based on Convolutional Neural Networks (CNNs) also have been developed [52]. Moreover, STraTS [53] first encodes the time series data into triple representations to avoid the limitations of using basic RNN and CNN in modeling irregular and sparse time series data and then builds the transformer-based forecasting model for modeling multivariate medical clinical time series. Graph-based time series forecasting methods can also be used as a self-supervised pretext task. Compared with RNNs and CNNs, Graph Neural Networks (GNNs) can better capture the correlation between variables in multivariate time series data, such as GDN [54].
Unlike the above methods, SSTSC [55] proposes a temporal relation learning prediction task based on the “Past-Anchor-Future” strategy as a self-supervised pretext task. Instead of directly forecasting the values of the future time windows, SSTSC predicts the relationships of the time windows, which can fully mine the temporal relationship structure in the data.
3.2 Autoencoder-based reconstruction
The autoencoder is an unsupervised artificial neural network composed of an encoder and a decoder [56]. The encoder maps the input $X$ to the representation $Z$, and then the decoder re-maps the representation $Z$ back to the input. The output of the decoder is defined as the reconstructed input $\tilde{X}$. The process can be expressed as:
$Z=E(X),\quad\tilde{X}=D(Z),$ (4)
where $E(\cdot)$ and $D(\cdot)$ represent the encoder and decoder, respectively. The difference between the original input $X$ and the reconstructed input $\tilde{X}$ is called the reconstruction error, and the goal of the self-supervised pretext task using autoencoder structure is to minimize the error between $X$ and $\tilde{X}$, i.e.,
$\mathcal{L}=\left\|X-\tilde{X}\right\|_{2}.$ (5)
The model structure of (4) is defined as the basic autoencoder (BAE). Most BAE-based methods jointly train the encoder $E(\cdot)$ and the decoder $D(\cdot)$. Then removing the decoder $D(\cdot)$ and leaving only the encoder $E(\cdot)$ that is used as a feature extractor, and the representation $Z$ is used for downstream tasks [57, 58, 59, 60]. For example, TimeNet [58], PT-LSTM-SAE[59], and Autowarp [57] all use RNN to build a sequence autoencoder model including encoder and decoder, which tries to reconstruct the input series. Once the model is learned, the encoder is used as a feature extractor to obtain an embedded representation of time series samples, which can help downstream tasks, such as classification and forecasting, achieve better performance. Zhang et al. [61] build a CNN-based autoencoder model and keep the encoder as a feature extractor after minimizing (5). The experimental results show that using the encoded representation is better than directly using the original time series data in industrial fault detection tasks.
However, the representations obtained by (5) are sometimes task-agnostic. Therefore, it is feasible to introduce additional training constraints based on (5). Abdulaal et al. [60] focus on the complex asynchronous multivariate time series data and introduce the spectral analysis in the autoencoder model. The synchronous representation of the time series is extracted by learning the phase information in the data, which is eventually used for the anomaly detection task. DTCR [62] is a temporal clustering-friendly representation learning model. It introduces K-means constraints in the reconstruction task, making the learned representation more friendly to clustering tasks. USAD [63] uses an encoder and two decoders to build an autoencoder model and introduces adversarial training based on (5) to enhance the representation ability of the model. FuSAGNet [64] introduces graph learning on the sparse autoencoder to model relationships in multivariate time series explicitly.
Denoising autoencoder (DAE) is another widely used approach, which is based on the addition of noise to the input series to corrupt the data, and then followed by the reconstruction task [65]. DAE can be formulated as:
$X_{n}=\mathcal{T}(X),\quad Z=E(X_{n}),\quad\tilde{X}=D(Z),$ (6)
where $\mathcal{T}$ indicates the operation that adds noise. The learning objective of a DAE is the same as that of a BAE, which is to minimize the difference between $X$ and $\tilde{X}$. In time series modeling, more than one method can add noise to the input, such as adding Gaussian noise [66, 67] and randomly setting some time steps to zero [68, 69].
(a) Autoregressive-based forecasting task
(b) Autoencoder-based reconstruction task
(c) Diffusion-based generation task
Figure 3: Three categories of generative-based SSL for time series data.
Mask autoencoder (MAE) is a structure widely used in language models and vision models in recent years [11, 70]. The core idea behind MAE is that in the pre-training phase, the model first masks part of the input and then tries to predict the masked part through the unmasked part. Unlike BAE and DAE, the loss of MAE is only computed on the masked part. MAE can be formulated as:
$X_{m}=\mathcal{M}(X),\quad Z=E(X_{m}),\quad\tilde{X}=D(Z),$ (7)
$\mathcal{L}=\mathcal{M}(\left\|X-\tilde{X}\right\|_{2}),$ (8)
where $\mathcal{M}(\cdot)$ represents the mask operation, and $X_{m}$ represents the masked input. In language models, since the input is usually a sentence, the mask operation masks some words in a sentence or replaces them with other words. In vision models, the mask operation will mask the pixels or patches in an image. For time series data, a feasible operation is to mask part of the time steps and then use the unmasked part to predict the masked time steps. Existing masking methods for time series data can be divided into three categories: time-step-wise masking, segment-wise masking, and variable-wise masking.
The time-step-wise masking randomly selects a certain proportion of time-steps in the series to mask, so the fine-grained information is easier to capture, but it is difficult to learn contextual semantic information in time series. The segment-wise masking randomly selects segments to mask, which allows the model to pay more attention to slow features in the time series, such as trends or high-level semantic information. STEP [71] divided the series into multiple non-overlapping segments of equal length and then randomly selected a certain proportion of the segments for masking. Moreover, STEP pointed out two advantages of using segment-wise masking: the ability to capture semantic information and reduce the input length to the encoder. Different from STEP, Zerveas et al. [72] performed a more complex masking operation on the time series, i.e., the multivariate time series was randomly divided into multiple non-overlapping segments of unequal length on each variable. Variable-wise masking was introduced by Chauhan et al.[73], who defined a new time series forecasting task called variable subset forecast (VSF). In VSF, the time series samples used for training and inference have different dimensions or variables, which may be caused by the absence of some sensor data. This new forecasting task brings the feasibility of self-supervised learning based on variable-wise masking. Unlike the random masking, TARNet [74] considers the pre-trained model based on the random masking strategy irrelevant to the downstream task, which leads to sub-optimal representations. Therefore, TARNet developed a new data-driven masking strategy that uses self-attention score distribution from downstream task training to determine the time steps associated with the downstream task and then masks out these necessary time steps and reconstructs them.
Variational autoencoder (VAE) is a model based on variational inference [75, 76]. The encoder encodes the input $X$ to the probability distribution $P(Z|X)$ instead of the explicit representation $Z$. When the decoder is used to reconstruct the input, a vector generated by sampling from the distribution $P(Z|X)$ will be used as input to the decoder. The process can be expressed as:
$P(Z|X)=E(X),\quad Z=\mathcal{S}(P(Z|X)),\quad\tilde{X}=D(Z),$ (9)
where $\mathcal{S}(\cdot)$ represents the sampling operation. Unlike (5), the loss function of a VAE includes two terms: the reconstruction item and the regularization item, i.e.,
$\mathcal{L}=\left\|X-\tilde{X}\right\|_{2}+\text{KL}(\mathcal{N}(\mu,\delta),\mathcal{N}(0,I)),$ (10)
where $\text{KL}(\cdot)$ represents the Kullback-Leibler divergence. The role of the regularization term is to ensure that the learned distribution $P(Z|X)$ is close to the standard normal distribution, thereby regulating the representation of the latent space. The representation learning method based on VAE can model the distribution of each time step to better capture the complex spatiotemporal dependencies and provide better interpretability in time series modeling tasks. For example, InterFusion [77] is a hierarchical VAE that models inter-variable and temporal dependencies in time series data. OmniAnomaly [78] combines VAE and Planar Normalizing Flow to propose an interpretable time series anomaly detection algorithm. In order to better capture the dependencies between different variables in multivariate time series, GRELEN [79] and VGCRN [80] introduce the graph structure and in VAE. In addition to modeling on regular time series, the methods based on VAE have made progress in sparse and irregular time series data representation learning, such as mTANs [81], HetVAE [82] and P-VAE [83]. The latest work attempts to extract seasonal and trend representations in time series data based on VAE. LaST [84] is a disentangled variational inference framework with mutual information constraints. It separates seasonal and trend representations in the latent space to achieve accurate time series forecasting.
3.3 Diffusion-based generation
As a new kind of deep generative model, diffusion models have achieved great success recently in many fields, including image synthesis, video generation, speech generation, bioinformatics, and natural language processing due to their powerful generating ability [85, 86, 87, 88, 89]. We provide a brief introduction to this class of methods and show the main applications in time series analysis recently.
The key design of the diffusion model contains two inverse processes: the forward process of injecting random noise to destruct data and the reverse process of sample generation from noise distribution (usually normal distribution). The intuition is that if the forward process is done step-by-step with a transition kernel between any two adjacent states, then the reverse process can follow a reverse state transition operation to generate samples from noise (the final state of the forward process). However, it is usually not easy to formulate the reverse transition kernel, and thus diffusion models learn to approximate the kernel by deep neural networks. Nowadays, there are mainly three basic formulations of diffusion models: denoising diffusion probabilistic models (DDPMs) [85, 90], score matching diffusion models [91, 92], and score SDEs [93, 94].
3.3.1 DDPMs
The forward and reverse processes in DDPMs are two Markov chains: a forward chain that adds random noise to data and a reverse chain that transforms noise back into data. Formally, denote the data distribution as $\boldsymbol{x}_{0}\sim q(\boldsymbol{x}_{0})$, the forward Markov process gradually adds Gaussian noise to the data according to transition kernel $q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$. It generates a sequence of random variables $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$. Thus the joint distribution of $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$ conditioned on $\boldsymbol{x}_{0}$ is
$q(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}|\boldsymbol{x}_{0})=\prod_{t=1}^{T}q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1}).$ (11)
For simplicity of calculation, the transition kernel is usually set as
$q(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})=\mathcal{N}(\boldsymbol{x}_{t};\sqrt{1-\beta_{t}}\boldsymbol{x}_{t-1},\beta_{t}I),$ (12)
where $\beta_{1},\beta_{2},\dots,\beta_{T}$ is a variance schedule of the forward process (usually chosen $\beta_{t}\in(0,1)$ ahead of model training) and $p(\boldsymbol{x}_{T})=\mathcal{N}(\boldsymbol{x}_{T};0,I).$ Similarly, the joint distribution of the reverse process is
$p_{\theta}(\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T})=p(\boldsymbol{x}_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (13)
where $\theta$ is the model parameters and $p_{\theta}(\boldsymbol{x_{t-1}|x_{t}})=\mathcal{N}(\boldsymbol{x}_{t-1};\mu_{\theta}(\boldsymbol{x}_{t},t),\sum_{\theta}(\boldsymbol{x}_{t},t)).$ The key to achieving the success of sample generating is training the parameters $\theta$ to match the actual reverse process, that is, minimizing the Kullback-Leibler divergence between the two joint distributions. Thus, the training loss is
$\textbf{KL}(q(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T})||p_{\theta}(\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T}))$ (14)
$\geq\mathbf{E}[-\log p_{\theta}(\boldsymbol{x}_{0})]+const,$
according to Jensen’s inequality. See more details in [87].
3.3.2 Score matching diffusion models
For a given probability density function $p(\boldsymbol{x})$, the Stein score (noted as $\nabla_{x}\log p(\boldsymbol{x})$), which is the gradient of the log-density function of the input data $\boldsymbol{x}$, points to the directions with the largest growth rate in probability density function. The key idea of score based diffusion models is to perturb data with a sequence of Gaussian noise and then jointly estimate the score functions for all noisy data distributions by training a deep neural network conditioned on noise levels. The motivation of the idea is that, in many situations, it is easier to model and estimate the score function than the original probability density function.
Langevin dynamics is one of the proper techniques. With a step size $\alpha>0$, a number of iterations $T$ and an initial sample $x_{0}$, Langevin dynamics iteratively does the following estimation to gain a close approximation of $p(\boldsymbol{x})$
$\boldsymbol{x}_{t}\leftarrow\boldsymbol{x}_{t-1}+\alpha\nabla_{x}\log p(\boldsymbol{x}_{t-1})+\sqrt{2\alpha}\boldsymbol{z}_{t},1\leq t\leq T,$ (15)
where $\boldsymbol{z}_{t}\sim\mathcal{N}(0,I)$. However, the score function is inaccurate without the training data, and Langevin dynamics may not converge correctly. Thus, the key approach (NCSN, a noise-conditional score network), perturbing data with a noise sequence and jointly estimating the score function for all the noisy data with a deep neural network conditioned on noise levels, is proposed [91]. Training and sampling are decoupled in score-based generative models, which inspires different choices in such two processes [92].
3.3.3 Score SDEs
DDPMs and SGMs are both limited to discrete and finite time steps (which can also be called noise levels). Score SDEs process the diffusion operation according to the stochastic differential equation (SDE) [94]:
$d\boldsymbol{x}=f(\boldsymbol{x},t)dt+g(t)d\boldsymbol{w},$ (16)
where $f(\boldsymbol{x},t)$ and $g(t)$ are diffusion function and drift function of the SDE, respectively, and $\boldsymbol{w}$ is a standard Wiener process. Different from DDPMs and SGMs, Score SDEs generalize the diffusion process to the case of infinite time steps. Fortunately, DDPMs and SGMs also can be formulated with corresponding SDEs. For DDPMs, the SDE is
$d\boldsymbol{x}=-\frac{1}{2}\beta(t)\boldsymbol{x}dt+\sqrt{\beta(t)}d\boldsymbol{w},$ (17)
where $\beta(\frac{t}{T})=T\beta_{t}$ when T goes to infinity; for SGMs, the SDE is
$d\boldsymbol{x}=\sqrt{\frac{d[\delta(t)^{2}]}{dt}}d\boldsymbol{w},$ (18)
where $\delta(\frac{t}{T})=\delta_{t}$ as T goes to infinity.
With any diffusion process in the form of (16), the reverse process can be gained by solving the following SDE:
$d\boldsymbol{x}=[f(\boldsymbol{x},t)-g(t)^{2}\nabla_{\boldsymbol{x}}\log q_{t}(\boldsymbol{x})]dt+g(t)d{\boldsymbol{\overline{w}}},$ (19)
where $\boldsymbol{\overline{w}}$ is a standard Wiener process when time flows reversely and $dt$ is an infinitesimal time step. Besides that, the existence of an ordinary differential equation, which is also called the probability flow ODE, is defined as follows.
$d\boldsymbol{x}=[f(\boldsymbol{x},t)-{\frac{1}{2}}{g(t)}^{2}\nabla_{\boldsymbol{x}}\log q_{t}(\boldsymbol{x})]dt.$ (20)
The trajectories of the probability flow ODE have the same marginals as the reverse-time SDE. Once the score function at each time step is known, the reverse SDE can be solved with various numerical techniques. Similar objective is designed with SGMs.
Diffusion models have also been applied in time series analysis recently. We will briefly introduce some of them, including the designed architectures and the main diffusion techniques used.
3.3.4 Diffusion models for time series imputation
Conditional Score-based Diffusion Models for Imputation (CSDI) [95] proposes a novel time series imputation method that utilizes score-based diffusion models conditioned on observed data. It is trained for imputation and can be easily generalized to forecasting by imputing data at the end of the time series sequence. To deal with the issue that the ground-truth missing values are not accessible in practice, a self-supervised training procedure is also proposed where values are masked as missing values to train the whole architecture.
3.3.5 Diffusion models for time series forecasting
Several works have appeared that take advantage of diffusion models to forecast time series. TimeGrad [96] takes an RNN conditioned diffusion probabilistic model at some time step to depict the fixed forward process and the learned reverse process. The cross-correlation of the time series is also depicted during the process stages. It is an autoregressive model for multivariate time series forecasting and learns gradients by optimizing the parameters that minimize the negative log-likelihood of the model:
$\sum_{t=t_{0}}^{T}-\log p_{\theta}(x_{t}^{0}|\boldsymbol{h}_{t-1}),$ (21)
[…]
4 Contrastive-based Methods
Contrastive learning is a widely used self-supervised learning strategy, showing a strong learning ability in computer vision and natural language processing. Unlike discriminative models that learn a mapping rule to true labels and generative models that try to reconstruct inputs, contrastive-based methods aim to learn data representations by contrasting between positive and negative samples. Specifically, positive samples should have similar representations, while negative samples have different representations. Therefore, the selection of positive samples and negative samples is very important to contrastive-based methods. This section sorts out and summarizes the existing contrastive-based methods in time series modeling according to the selection of positive and negative samples.
(a) Sampling contrast
(b) Prediction contrast
(c) Augmentation contrast
(d) Prototype contrast
(e) Expert knowledge contrast
Figure 4: Five categories of contrastive-based SSL for time series data.
4.1 Sampling contrast
Sampling contrast follows a widely used assumption in time series analysis that two neighboring time windows or time stamps have a high degree of similarity, so positive and negative samples are directly sampled from the raw time series, as shown in Fig. 4(a). Specifically, given a time window (or a timestamp) as an anchor, its nearby window (or the time stamp) is more likely to be similar (small distance), and the distant window (or the time stamp) should be less similar (large distance). The term “similar” indicates that two windows (or two-time stamps) have more common patterns, such as the same amplitude, the same periodicity, and the same trend.
As mentioned in [101], suppose one anchor $x^{ref}$, one positive sample $x^{pos}$, and $K$ negative samples ${x^{neg}_{k}}_{,\ k\in 1,2,\cdots,K}$ are chosen, we expect to assimilate $x^{ref}$ and $x^{pos}$ and to distinguish between $x^{ref}$ and $x^{neg}_{k}$, i.e.,
$\mathcal{L}=-\log(S(x^{ref},x^{pos}))-\sum_{k=1}^{K}\log(-S(x^{ref},x^{neg}_{k})),$ (22)
where $S(\cdot)$ denotes the similarity of the two representations. However, due to the non-stationary characteristics of most time series data, it is still a challenge to choose the correct positive and negative samples based on contextual information in time series data. Temporal neighborhood coding (TNC) was recently proposed to deal with this problem [102]. TNC uses Augmented Dickey-Fuller (ADF) statistical test to determine the stationary region and introduces PU (Positive-Unlabeled) learning to handle the problem of sampling bias by treating negative samples as unknown samples and then assigning weights to these samples. The learning objective is extended to
$\mathcal{L}=$ $-\mathbb{E}_{x^{pos}[\in\mathcal{N}}\log S(x^{ref},x^{pos})]$ (23)
$-\mathbb{E}_{x^{neg}\in\tilde{\mathcal{N}}}[(1-w)\times\log-S(x^{ref},x^{neg})$
$+w\times\log S(x^{ref},x^{neg})],$
where $w$ is the probability of sampling false negative samples. $\mathcal{N}$ denotes the neighboring area and $\tilde{\mathcal{N}}$ denotes the non-neighboring area. Supervised contrastive learning (SCL) [103] effectively addresses sampling bias, so introducing the supervised signal to identify positive and negative samples is a feasible solution. Neighborhood contrastive learning (NCL) is a recent time series modeling method that combines context sampling and the supervised signal to generate positive and negative samples [104]. NCL assumes that if two samples share some predefined attributes, then they are considered to share the same neighboring area.
4.2 Prediction contrast
In this category, prediction tasks that use the context (present) to predict the target (future information) are considered self-supervised pretext tasks, and the goal is to maximally preserve the mutual information of the context and the target. Contrastive predictive coding (CPC) proposed by [22] provides a contrastive learning framework to perform the prediction task using InfoNCE loss. As shown in Figure 4(b), the context $c_{t}$ and the sample from $p(x_{t+k}|c_{t})$ constructs positive pairs, and the samples form the ‘proposal’ distribution $p(x_{t+k})$ are negative samples. The learning objective is as follows:
[…]
where $f_{k}(\cdot)$ is the density ratio that preserves the mutual information of $c_{t}$ and $x_{t+k}$ [22], and it can be estimated by a simple log-bilinear model:
$f_{k}\left(x_{t+k},c_{t}\right)=\exp(z^{T}_{t+k}W_{k}c_{t}).$ (25)
[…]
In addition to the basic contextual prediction tasks mentioned before, some more complex prediction tasks were constructed and proved useful. CMLF [110] transforms time series into coarse-grained and fine-grained representations and proposes a multi-granularity prediction task. This allows the model to represent the time series at different scales. TS-TCC[111] and its extended version CA-TCC[112] designed a cross prediction task, which uses the context of $x^{T1}$ to predict the target in $x^{T2}$, and vice versa uses the context of $x^{T2}$ to predict the target in $x^{T1}$.
4.3 Augmentation contrast
Augmentation contrast is one of the most widely used contrastive frameworks, as shown in Fig. 4(c). Most methods utilize data augmentation techniques to generate different views of an input sample and then learn representations by maximizing the similarity of the views that come from the same sample and minimizing the similarity of the views that come from the different samples. SimCLR [16] is a very typical multi-view invariance-based representation learning framework, which has been used in many subsequent methods. The objective function based on this framework is:
$\mathcal{L}=-\log\frac{\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{1},\boldsymbol{z}_{2}\right)/\tau\right)}{\sum_{k=1}^{2N}\mathbbm{1}_{[k\neq 1]}\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{1},\boldsymbol{z}_{k}\right)/\tau\right)},$ (26)
where $\tau$ is temperature parameter, $\operatorname{sim}(\cdot)$ represents the similarity between two representation vectors, and $z_{k}$ represents the training samples in a batch. It can be considered that in the feature learning framework based on multi-view invariance, the core is to obtain different views of the input samples. When handling images in computer vision, commonly used data augmentation methods include cropping, scaling, adding noise, rotation, and resizing [16]. However, compared with augmentation methods for images, the augmentation methods for time series needs to consider both temporal and variable dependencies.
Since time series data can be converted to frequency domain representations through Fourier transform, the augmentation method can be developed from the time and frequency domains. In the time domain, TS-TCC[111] and its extended version CA-TCC[112] designed two time series data augmentation techniques, one is strong augmentation (Permutation-and-Jitter), and the other is weak augmentation (Jitter-and-Scale). TS2Vec [113] generates different views through masking operations that randomly mask out some time steps. Generally speaking, there is no one-size-fits-all answer to the choice of data augmentation methods. Therefore, some works comprehensively compare and study the augmentation methods and further evaluate the performance on different tasks [114, 33, 115, 116].
Data augmentation in the frequency domain is also feasible for time series data. CoST [117] is a disentangled seasonal-trend representation learning method, which uses Fast Fourier Transform to convert different augmented views into amplitude and phase representations, and then uses (26) to train the model. BTSF [118] is a contrastive-based method based on a time-frequency fusion strategy, which first generates an augmented view in the time domain through the dropout operation and then generates another augmented view in the frequency domain through Fourier transform. Finally, the bilinear temporal-spectral fusion mechanism is used to achieve the fusion of time-frequency information. However, CoST and BTSF do not directly modify the frequency representation, while TF-C [119] is the first work to augment the time series data through frequency perturbations, which has achieved better performance than TS2Vec [113] and TS-TCC [111]. TF-C implements three augmentation strategies: Low- vs. High-band perturbations, Single- vs. Multi-component perturbations, and Random vs. Distributional perturbations.
In addition to the above methods, many view generation methods are closely related to downstream tasks. Recently, DCdetector [120] proposes a dual attention contrastive representation framework for time series anomaly detection. The in-patch and patch-wise representations are designed to gain two views of the input samples, as normal samples behave differently from abnormal ones in such two views. TimeCLR [121] proposed DTW augmentation, which can not only simulate phase shifts and amplitude changes but also retain the structure and characteristics of the time series. CLOCS [122] is a self-supervised pre-training method for medical and physiological signals, which uses multi-view invariance contrast in the three perspectives of time, space, and patient to promote higher similarity of representations from the same source. CLUDA [123] introduces multi-view invariance contrast in the time series domain adaptation problem, which captures the contextual representation of time series data through intra-domain and inter-domain contrast. MTFCC [124] is another view generation method based on multi-scale characteristics, which samples time series samples at multiple scales and considers that the views from the same sample have similar representations, even if their scales are different. Methods for constructing multiple contrastive views based on multi-granularity or multi-scale augmentations also include MRLF[125], CMLF [110], and SSLAPP[126].
4.4 Prototype contrast
The contrastive learning framework based on (24) and (26) is essentially an instance discrimination task, which encourages samples to form a uniform distribution in the feature space [127]. However, the real data distribution should satisfy that the samples of the same class are more concentrated in a cluster, while the distance between different clusters should be farther. SCL [103] is an ideal solution when real labels are available, but this is difficult to implement in practice, especially for time series data. Therefore, introducing clustering constraints into existing contrastive learning frameworks is an alternative, such as CC [128], PCL [129], and SwAV [36]. PCL and SwAV contrast the samples with the constructed prototypes, i.e., the cluster centers, which reduces the computation and encourages the samples to present a cluster-friendly distribution in the feature space. An illustration of prototype contrast is shown in Fig. 4(d).
In time series modeling based on prototypes contrast, ShapeNet [130] takes Shaplets as input and constructs a cluster-level triplet loss, which considers the distance between the anchor and multiple positive (negative) samples as well as the distance between positive (negative) samples. ShapeNet is an implicit prototype contrast because it does not introduce explicit prototypes (cluster centers) during the training phase. TapNet [131] and VSML [132] are explicit prototypes contrast because explicit prototypes are introduced. TapNet introduces a learnable prototype for each predefined class and classifies the input time series sample according to the distance between the sample and each class prototype. VSML defines virtual sequences, which have the same function as prototypes, i.e., minimize the distance between samples and virtual sequences, but maximize the distance between virtual sequences.
MHCCL [133] proposes a hierarchical clustering based on the upward masking strategy and a contrastive pairs selection strategy based on the downward masking strategy. In the upward mask strategy, MHCCL believes that outliers greatly impact prototypes, so these outliers should be removed when updating prototypes. The downward masking strategy, in turn, uses the clustering results to select positive and negative samples, i.e., samples belonging to the same prototype are regarded as true positive samples, and samples belonging to different prototypes are regarded as true negative samples.
4.5 Expert knowledge contrast
Expert knowledge contrast is a relatively new representation learning framework. Generally speaking, this modeling framework incorporates expert prior knowledge or information into deep neural networks to guide model training [134, 135]. In the contrastive learning framework, prior knowledge can help the model choose the correct positive and negative samples during training. An example of expert knowledge contrast is shown in Fig. 4(e). Here we sort out three interesting works.
Shi et al. [136] used the DTW distance of time series samples as prior information and believed that two samples with a small DTW distance have a higher similarity. Specifically, given the anchor $x^{ref}$ and the other two samples $x_{i}$ and $x_{j}$, the DTW distance between $x^{ref}$ and the other two samples is calculated first, then the sample with a small distance from $x^{ref}$ is considered as the positive sample of $x^{ref}$. This selection process is defined as
$\text{label}=\left\{\begin{array}[]{ll}1,&\mathrm{DTW}\left(\boldsymbol{x}^{ref},\boldsymbol{x}_{i}\right)\geq\operatorname{DTW}\left(\boldsymbol{x}^{ref},\boldsymbol{x}_{j}\right)\\ 0,&\text{ otherwise }\end{array}.\right.$ (27)
Based on Pair-Loss, ExpCLR[137] introduces expert features of time series data to obtain more informative representations. Given two input samples $x_{i}$ and $x_{j}$ and corresponding representations $f_{i}$ and $f_{j}$, ExpCLR defines the normalized distance between two samples:
$s_{ij}=1-\frac{\left\|f_{i}-f_{j}\right\|_{2}}{\mathtt{max}\left\|f_{k}-f_{l}\right\|}_{2},$ (28)
where $f_{k}$ and $f_{l}$ are the two representation vectors with the largest distance among all samples. Compared with the original Pair-Loss, the distance between samples $x_{i}$ and $x_{j}$ is changed from a discrete value (0 and 1) to a continuous value $s_{ij}$, which enables the model to learn more accurately about the relationship between samples, thus thereby enhancing the representation ability of the model. In addition to the above two works, SleepPriorCL[138] was proposed to alleviate the sampling bias problem faced by (26). Like ExpCLR, SleepPriorCL also introduces prior features to ensure the model can identify correct positive and negative samples.
5 Adversarial-based Methods
Adversarial-based self-supervised representation learning methods utilize generative adversarial networks (GANs) to construct pretext tasks. GAN contains a generator $\mathcal{G}$ and a discriminator $\mathcal{D}$. The generator $\mathcal{G}$ is responsible for generating synthetic data similar to real data, while the discriminator $\mathcal{D}$ is responsible for determining whether the generated data is real data or synthetic data. Therefore, the goal of the generator is to maximize the decision failure rate of the discriminator, and the goal of the discriminator is to minimize its failure rate [139, 48]. The generator $\mathcal{G}$ and the discriminator $\mathcal{D}$ are a mutual game relationship, so the learning objective is:
$\mathcal{L}=\mathbb{E}_{x\sim\mathcal{P}_{\text{data}}(x)}[\log\mathcal{D}(x)]+\mathbb{E}_{z\sim\mathcal{P}_{\mathbf{z}}(z)}[\log(1-\mathcal{D}(\mathcal{G}(\mathbf{z})))].$ (29)
According to the final task, the existing adversarial-based representation learning methods can be divided into time series generation and imputation, and auxiliary representation enhancement. The examples of these adversarial-based SSL are shown in Fig. 5.
(a) time series generation
(b) time series imputation
(c) Auxiliary representation enhancement
Figure 5: Three categories of adversarial-based SSL for time series data.
5.1 Time series generation and imputation
[…]
Complete time series generation refers to generating a new time series that does not exist in the existing data set. The new sample can be a univariate or multivariate time series. C-RNN-GAN [145] is an early method of generating time series samples using GAN. The generator is an RNN, and the discriminator is a bidirectional RNN. RNN-based structures can capture the dynamic dependencies in multiple time steps but ignore the static features of the data. TimeGAN [146] is an improved time series generation framework that combines the basic GAN with the autoregressive model, allowing the preservation of temporal dynamic characteristics of the series. TimeGAN also emphasizes that static features and temporal characteristics are crucial to the generation task.
Some recently proposed methods consider more complex time series generative tasks[149, 150, 147, 151]. For example, COSCI-GAN [149] is a time series generation framework that considers the correlation between each dimension of the multivariate time series. It includes Channel GANs and Central Discriminator. Channel GANs are responsible for generating data in each dimension independently, while Central Discriminator is responsible for determining whether the correlation between different dimensions of the generated series is the same as the raw series. PSA-GAN [150] is a framework for long-time series generation and introduces a self-attention mechanism. It further presents Context-FID, a new metric for evaluating the quality of generated series. Li et al. [147] explored the generation of time series data with irregular spatiotemporal relationships and proposed TTS-GAN, which uses a Transformer instead of an RNN to build the discriminator and the generator and treats the time series data as image data of height one.
[…]
5.2 Auxiliary representation enhancement
In addition to generation and interpolation tasks, an adversarial-based representation learning strategy can be added to existing learning frameworks as additional auxiliary learning modules, which we call adversarial-based auxiliary representation enhancement. The auxiliary representation enhancement aims to promote the model to learn more informative representations for downstream tasks by adding adversarial-based learning strategies. It can be defined as:
$\mathcal{L}=\mathcal{L}_{base}+\mathcal{L}_{adv},$ (30)
where $\mathcal{L}_{base}$ is the basic learning objective and $\mathcal{L}_{adv}$ is the additional adversarial-based learning objective. It should be noted that when $\mathcal{L}_{adv}$ is not available, the model can still extract representations from the data, so $\mathcal{L}_{adv}$ is regarded as an auxiliary learning objective.
USAD [63] is a time series anomaly detection framework that includes two BAE models, and two BAE are defined as $AE_{1}$ and $AE_{2}$, respectively. The core idea behind USAD is to amplify the reconstruction error by adversarial training between two BAEs. In USAD, $AE_{1}$ is regarded as the generator, and $AE_{2}$ is regarded as the discriminator. The auxiliary goal is to use $AE_{2}$ to distinguish real data from reconstructed data from $AE_{1}$, and train $AE_{1}$ to deceive $AE_{2}$, the whole process can be expressed as:
$\mathcal{L}_{adv}=\min_{AE_{1}}\max_{AE_{2}}\left\|{W-AE_{2}(AE_{1}(W))}\right\|_{2},$ (31)
where $W$ is the real input series. Similar to USAD, AnomalyTrans [154] also uses an adversarial strategy to amplify the anomaly score of anomalies. But unlike (31), which uses reconstruction error, AnomalyTrans defines prior-association and series-association and then uses the Kulback-Leibler divergence to measure the error of the two associations.
DUBCNs [155] and CRLI [156] are used for series retrieval and clustering tasks, respectively. Both methods adopt RNN-based BAE as the model, and the clustering-based loss and adversarial-based loss are added to the basic reconstruction loss, i.e.,
$\mathcal{L}=\mathcal{L}_{mse}+\lambda_{1}\mathcal{L}_{cluster}+\lambda_{2}\mathcal{L}_{adv}.$ (32)
where $\lambda_{1}$ and $\lambda_{2}$ are the weight coefficients of the auxiliary objective.
The adversarial-based strategy is also effective in other time series modeling tasks. For example, introducing adversarial training in time series forecasting can improve the accuracy and capture long-term repeated patterns, such as AST [157] and ACT [158]. BeatGAN [159] introduces adversarial representation learning in the abnormal beat detection task of ECG data and provides an interpretable detection framework. In modeling behavior data, Activity2vec [160] uses adversarial-based training to model target invariance and enhance the representation ability of the model in different behavior stages.
6 Applications and Datasets
TABLE II: Summary of time series applications and widely used datasets. The UCR and UEA datasets have multiple sub-datasets respectively, so their size and dimension are not fixed values. The size and dimension of each sub-dataset are represented by $\mathcal{M}$ and $\mathcal{D}$, respectively. AnRa represents Anomaly Ratio. SaIn represents Sampling Interval.
Applications Dataset Size Dimension Source Citations Comment
[…]
SSL has many applications across different time series tasks, such as forecasting, classification, anomaly detection, etc. We summarize the most widely used datasets and representative references according to the application area. In this article, we collect widely used datasets in four tasks: anomaly detection, forecasting, classification, and clustering. As shown in Table II, we provide useful information, including dataset name, dimension, size, source, citations, and useful comments. Datasets for classification tasks and clustering tasks are listed together because the goals of these two tasks are consistent.
Anomaly Detection. The anomaly detection problem for time series is usually formulated as identifying outlier time points or unexpected time sequences relative to some norm or usual signal. Most time series anomaly detection methods are constructed under an unsupervised learning framework because obtaining labels for anomalous data is challenging. Autoregressive-based forecasting and autoencoder-based reconstruction are the most commonly used modeling strategies. To be concrete, THOC [51] and GDN [54] employ autoregressive-based forecasting SSL framework, which assumes that anomalous sequences or time points are not predictable. While based on the principle that anomalous are hard to reconstruct, RANSynCoders [60], USAD [63], AnomalyTrans [154], and DAEMON [165] employ autoencoder-based reconstruction SSL framework. Furthermore, VGCRN [80] and FuSAGNet [64] combine two frameworks to achieve more robust and accurate detection results. It is beneficial to introduce an adversarial-based SSL method in the detection models, which can further amplify the difference between normal and anomalous data, such as USAD [63] and DAEMON [165].
Forecasting. Time series forecasting is the process of analyzing time series data using statistics and modeling to make predictions of future windows or time points. The pretext task based on autoregressive-based forecasting is essentially a time series forecasting task. Therefore, various models based on forecasting tasks are proposed, such as Pyraformer [174], FilM [15], Quatformer [175], Informer [171], Triformer [176], Scaleformer[187], Crossformer [189], and Timesnet [172]. Moreover, we found that decomposing the series (seasonality and trend) and then learning and forecasting on the decomposed components will help improve the final forecasting accuracy, such as MICN [184] and CoST [117]. Besides, introducing an adversarial SSL is viable when missing values are in the series. For example, LGnet [200] introduces adversarial training to enhance the modeling of global temporal distribution, which mitigates the impact of missing values on forecasting accuracy.
Classification and Clustering. The goal of classification and clustering tasks is the same, i.e., to identify the real category to which a certain time series sample belongs. Contrastive-based SSL methods are the most suitable choice for these two tasks since the core of contrastive learning is identifying positive and negative samples. Specifically, TS-TCC [111] introduces temporal contrast and contextual contrast in order to obtain more robust representations. TS2Vec [113] and MHCCL [133] performs a hierarchical contrastive learning strategy over augmented views, which enables robust representations. Similar to anomaly detection and prediction tasks, an adversarial-based SSL strategy can also be introduced into classification and clustering tasks. DTCR [62] propose a fake-sample generation strategy to assist the encoder in obtaining more expressive representations.
In summary, generative-based SSL is more suitable for anomaly detection and forecasting tasks, while contrastive-based SSL is more suitable for classification and clustering tasks. Adversarial-based SSL can play a role in various tasks, but in most cases, it is used as an additional regularization term to ensure that the features extracted by the model are more robust and informative. The above results and work also show that a hybrid approach (mixing multiple SSL strategies) is a better choice.
7 Discussion and Future Directions
Though promising progress has been made, the development of time series SSL is still at a developing stage. In this section, we point out some critical problems in current studies and outline several research directions worthy of further investigation.
7.1 Selection and combination of data augmentation
Data augmentation is one of the effective methods to generate augmented views in SSCL [201, 46]. The widely used methods for time series data include jitter, scaling, rotation, permutation, and warping [114, 33, 115, 116, 202]. In SimCLR [16], nine different augmentation methods for image data were discussed. The experiments show that “no single transformation suffices to learn good representations” and “the composition of random cropping and random color distortion is the most effective augmentation method”. This naturally raises the question of which one or a composition of data augmentation methods is optimal for time series SSCL. Recently, Um et al. [203] show that the combination of three basic augmentation methods (permutation, rotation, and time warping) is better than that of a single method and achieves the best performance in time series classification task. Iwana et al. [116] evaluate twelve time series data augmentation methods on 128 time series classification datasets with six different types of neural networks. The results show that the window warping and the window slicing algorithms proposed by Le et al. [204] are the most recommended algorithms. Pöppelbaum et al. [114] compared ten augmentation methods in time series classification tasks. The results show that left-to-right flipping, crop-and-resize, time warping, random smoothing, and noise appear to be the frontrunner approaches. Different evaluation frameworks give different recommendations and results. Therefore, an interesting direction is to construct a reasonable evaluation framework for time series data augmentation methods, then further select the optimal method or combination strategy.
7.2 Inductive bias for time series SSL
Existing SSL methods often pursue an entirely data-driven modeling approach. However, introducing reasonable inductive bias or prior is helpful for many deep neural networks-based modeling tasks [134, 205, 206]. On the one hand, although a purely data-driven model can be easily extended to various tasks, it requires much data to train it. On the other hand, time series data usually has some available characteristics, such as seasonal, periodic, trend, and frequency domain biases [207, 208, 209]. Thus one future direction is to consider more effective ways to induce inductive biases into time series SSL based on the understanding of time series data and characteristics of specific tasks.
7.3 SSL for irregular and sparse time series
Irregular and sparse time series data also widely exist in various scenarios. This data is measured at irregular time intervals, and not all the variables are available for each sample [210]. The straightforward approach to deal with irregular and sparse time series data is to use interpolation algorithms to estimate missing values [152, 153, 211]. However, interpolation-based models add undesirable noise and extra overhead to the model which usually worsens as the time series become increasingly sparse [53]. Moreover, irregular and sparse time series data is often expensive to obtain sufficient labeled data, which motivates us to build time series analysis models based on SSL in various tasks. Therefore, building SSL models directly on irregular and sparse time series data without interpolation is a valuable direction.
7.4 Theoretical analysis of contrastive learning
Contrastive learning pre-training on unlabeled data followed by supervised fine-tuning on labeled data is a widely used paradigm for learning from limited labeled data. Even the pre-training and downstream tasks have different training constraints. This has triggered many scholars to explore the theoretical basis behind SSCL. For example, Wang et al. [127] prove that the contrastive loss optimizes two key properties: alignment and uniformity, and introduce an optimizable metric to quantify each property. Wang et al. [212] build relations between the uniformity and the temperature $\tau$ in contrastive loss and point out that the temperature plays a key role in controlling the local separation and global uniformity of the embedding. Another strategy starts from the perspective of data augmentation methods and augmented views [213, 214, 215, 216]. Both Zhao et al. [214] and Wang et al. [216] point out that the transferability of contrastive learning depends mainly on the choice of data augmentation, and the learning process is to gradually cluster intra-class samples by aligning positive samples. Moreover, Tian et al. [213] start by analyzing the mutual information between augmented views and argue that the mutual information between views should be reduced while the relevance of augmented views to downstream tasks needs to be guaranteed. The existing methods provide an understanding and theoretical analysis of contrastive learning from multiple perspectives. However, there are still challenges in this area, such as the theoretical understanding of free negatives contrastive learning.
7.5 Pretraining and large models for time series
Nowadays, many large language models have shown their powerful perception and learning capability for many different tasks. A similar phenomenon also appears in computational vision [217]. It is naturally an interesting question, how about the time series analysis field? As far as we know, there is limited work on pretraining models in large-scale time series. Exploring the potentiality of pretraining and large time series models is a promising direction.
7.6 Adversarial attacks and robust analysis on time series
With the widespread use of deep neural networks in time series forecasting, classification, and anomaly detection, the vulnerability and robustness of deep models under adversarial have become a significant concern [218]. In the field of time series forecasting, Liu et al. [219] study the indirect and sparse adversarial attacks on multivariate probabilistic forecasting models for time series forecasting and propose two defense mechanisms: randomized smoothing and mini-max defense. Wu et al. [220] propose an attack strategy for generating an adversarial time series by adding malicious perturbations to the original time series to deteriorate the performance of time series prediction models. For time series classification, Karim et al. [221] focus on non-differentiable models (such as 1-NN DTW) in the time series domain and propose the student-teacher framework as a general solution to perform a proxy attack on the target model, differentiable or not. Zhuo et al. [222] summarize and compare various recent and typical adversarial attack and defense methods for fault classifiers in data-driven fault detection and classification systems, including white-box attack (FGSM[223], IGSM[224], C&\&&W attack[225], DeepFool[226]) and gray-box and black-box attack (UAP [227], SPSA [228], Random noise). The research on adversarial attacks and defenses against time series data is a worthwhile direction, but there is much less literature on this topic. Existing studies mainly involve forecasting and classification tasks. However, the impact of adversarial examples on time series self-supervised pre-training tasks is still unknown.
7.7 Time series SSL in collaborative systems
Distributed systems have been widely deployed in many scenarios, including intelligent control systems, wireless sensor networks, network file systems, the Internet of Things, etc. On the one hand, an appropriate collaborative learning strategy is fundamental in these systems, as users can train their own local models without sharing their private local data and circumvent the relevant privacy policy [229]. On the other hand, time series data is also widely distributed in various places in the system, and obtaining sufficient labeled data is also difficult, so time series SSL has great deployment potential. In recent years, federated learning has been the most popular collaborative learning framework and has been used successfully in various applications. Combining time series self-supervised learning and federated learning is a valuable research direction that can provide additional modeling tools for modern distributed systems.
7.8 Benchmark evaluation for time series SSL
SSL has many applications in time series classification, forecasting, clustering, and anomaly detection. However, most current research seeks to achieve the best performance on specific tasks and needs more discussion and evaluation of the self-supervised technique. Another interesting direction is to pay more attention to SSL, analyze its properties in time series modeling tasks, and give reliable benchmark evaluation.
8 Conclusion
This article concentrates on time series SSL methods and provides a new taxonomy. We categorize the existing methods into three broad categories according to their learning paradigms: generative-based, contrastive-based, and adversarial-based. Moreover, we sort out all methods into ten detailed subcategories: autoregressive-based forecasting, autoencoder-based reconstruction, diffusion-based generation, sampling contrast, prediction contrast, augmentation contrast, prototypes contrast, expert knowledge contrast, generation and imputation, and auxiliary representation enhancement. We also provide useful information about applications and widely used time series datasets. Finally, multiple future directions are summarized. We believe this review fills the gap in time series SSL and ignites further research interests in SSL for time series data.